OpenStudy (anonymous):

a coil of a wire is situated in a 0.5 T uniform magnetic field. The area of a coil is 2.0m^2.what is the magnetic flux if the angle between the magnetic field and the normal to the surface of the soil is 60 degrees ? after 5s, the magnetic field is now parallel to the normal to the surface. what is the induced emf ?? .. PLEASE .. I NEED THE ANSWERS .. :)) HELP MEEEEE

OpenStudy (anonymous):

flux linked is given by\[\phi = E S\cos \theta\] here area is taken as a vector in the direction of normal hebce in the given case phi= 05*2*1/2=1/2H/m^2 Induced emf is \[\epsilon= d \phi /dt\] here change flux is 1/2 H/m^2 hence induced emf is 1/2*1/5=0.1V